Grade 12Algebra

Vidya and Pratap went for a picnic. Their mother gave them a water bottle that contained 5 liters of water. Vidya consumed 2/5 of the water. Pratap consumed the remaining water.
What fraction of the total quantity of water did Pratap drink?

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

From the question, it is given that,
Amount of water in the water bottle = 5 liters
Then,
Amount of water consumed by Pratap = (1–water consumed by Vidya)
= (1–(2/5))
= (5-2)/5
= 3/5
∴Total amount of water consumed by Pratap = 3/5 of 5 liters
= (3/5) × 5
= 3 liters
So, the total amount of water drank by Pratap is 3 liters

Thanks